OpenStudy (anonymous):

OMG SOMEONE HELP ME FOREAL The average daily air temperature (F) for Fairbanks, Alaska, from 1941 to 1970, can be modeled by the equation: t(x)= 37sin[2/365(x-100)] + 25 where x represents the first of January. On what dates do you expect the average temperature to be 32 F ? 32= 37sin [2pi/365(x-100)] +25 I don't know how To solve for x

OpenStudy (ajanijones):

:lol

OpenStudy (ajanijones):

OMG SOMEONE HELP ME FOREAL

OpenStudy (anonymous):

lol ... i've bumped this like 4 times

OpenStudy (ajanijones):

I'll try and help see what i can do.

OpenStudy (anonymous):

Gracias :)

OpenStudy (turingtest):

start solving it like you would a regular algebra problem until you get the sine term by itself on one side of the equation, and a constant on the other 32= 37sin [2pi/365(x-100)] +25 if this is hurting your eyes, first rewrite it as 32=37u+25 where u=sin [2pi/365(x-100)] and solve for u, then dell me what you get

OpenStudy (turingtest):

tell*

OpenStudy (anonymous):

Aghh okay give me a minute

OpenStudy (anonymous):

well to be honest i don't know where to start. I really don't have the brains

OpenStudy (turingtest):

if it were 32=37u+25 could you solve for u?

OpenStudy (anonymous):

i put the u on the left side, and 32 on the right side?

OpenStudy (anonymous):

to get u by itself? idk

OpenStudy (anonymous):

or i subtract 25 from both sides?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

that is step 1

OpenStudy (anonymous):

so 7= 37u ?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

than i divide?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

would it be 37u/7 ?

OpenStudy (turingtest):

you want u by itself, so no

OpenStudy (anonymous):

than how do i divide..

OpenStudy (turingtest):

there are only two obvious choices: divide both sides by 7 or divide both sides by 37. only one gives u a coefficient of 1....

OpenStudy (anonymous):

Okay yeah you divide, but i don't understand how you get u by itself if it's with 37? i really don't know

OpenStudy (anonymous):

7/37 = u?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

and then what

OpenStudy (turingtest):

now substitute back in u=sin [2pi/365(x-100)]

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

where does 7/37 go where x is?

OpenStudy (turingtest):

no, you put sin [2pi/365(x-100)] were the u is I really think trigonometry is going to be very difficult without more algebra skills.

OpenStudy (anonymous):

yeah i don't know anything.

OpenStudy (anonymous):

okay but idk what to do after that

OpenStudy (turingtest):

take the inverse sine of both sides

OpenStudy (anonymous):

i don't know what that means

OpenStudy (turingtest):

then you have not been studying properly. You can't answer this question without knowledge of the sine function and it's inverse

OpenStudy (anonymous):

i don't understand none of this.

OpenStudy (turingtest):

\[\sin x=u\iff x=\sin^{-1}u\]

OpenStudy (anonymous):

oh fk it. nvm later.

OpenStudy (turingtest):

that is how an inverse sine works. I suggest you go back and reread the beginning of your book or whatever material you have. This question is impossible without those tools at your disposal.